A rabbit run across a parking lot. The coordinate of the rabbits as a function of time t are
x=0.3t square +7.2t square+28
Y=0.22t square-9.1t+80 with the in seconds and x and you in meters
Find the velocity of rabbit at time t=15s
1
Expert's answer
2018-12-25T08:38:11-0500

Components of the velocity

v_x (t)=dx/dt=0.6t+7.2

v_y (t)=dy/dt=0.44t-9.1

At the time t=15s

v_x (15)=0.6×15+7.2=16.2 m/s

v_y (15)=0.44×15-9.1=-2.5 m/s

The absolute value of the rabbit velocity

v=√(v_x^2+v_y^2 )=√(〖16.2〗^2+(-2.5)^2 )=16.4 m/s